# thanks: Corresponding author

### v1

@@ IV Discussion
Figure 2: 5-year limit results of CTA with and without correction.
@@
Figure 3 presents a comparison between our constraint results and those obtained by other detectors based on observational data Linton et al. (2006); Alexandreas et al. (1993); Amenomori, M and Cao, Z and Dai, BZ and Ding, LK and Feng, YX and Feng, ZY and Hibino, K and Hotta, N and Huang, Q and Huo, AX and others (1995); Archambault (2018); Abdo et al. (2015); Ackermann et al. (2018); Albert et al. (2020); Aharonian et al. (2023); Lopez-Coto et al. (2022); Yang et al. (2024). It is evident that CTA exhibits significantly higher sensitivity in constraining the local PBH burst rate density compared to most previous detectors. The sensitivity of CTA is only slightly lower than that of LHAASO for bursts with durations shorter than 0.01 s and lower than that of the future SWGO detector for bursts with durations shorter than 1 s. This demonstrates that CTA holds a distinct advantage in searching for long-duration PBH bursts.
@@
Figure 3: The PBH burst local rate density upper limits of CTA at 99% confidence level, compared with the results from Whipple Linton et al. (2006), CYGNUS Alexandreas et al. (1993), H.E.S.S. Aharonian et al. (2023), Tibet Air Shower Array Amenomori, M and Cao, Z and Dai, BZ and Ding, LK and Feng, YX and Feng, ZY and Hibino, K and Hotta, N and Huang, Q and Huo, AX and others (1995), Milagro Abdo et al. (2015), VERITAS Archambault (2018), Fermi-LAT Ackermann et al. (2018), HAWC Albert et al. (2020), LHAASO Yang et al. (2024) and SWGO Lopez-Coto et al. (2022).
@@
To establish a reference for selecting the optimal burst duration when searching for PBH bursts in future observational studies, we analyze the burst durations to which CTA is most sensitive. This analysis is a key focus of our work. As shown in Table 1, better constraints are likely achieved by targeting PBHs with longer remaining lifetimes. However, since telescopes typically do not observe the same region for extended periods, PBHs with remaining lifetimes significantly longer than the observation time emit more photons than can be detected. Considering this limitation, we apply a rough correction for observations lasting longer than $10^{2}$ s to determine the optimal constraints on PBH bursts for CTA.
We extended the analysis to include remaining lifetimes up to $10^{5}$ s. For cases with remaining lifetimes of $10^{3}$ s and $10^{4}$ s, we assume that only 50% of the total photons emitted by a PBH are received during the observation period. For a remaining lifetime of $10^{5}$ s, additional factors must be considered, such as the inability to observe continuously through the night and variations in nighttime duration due to the detector latitude and the Earth’s revolution. The CTA south array is located at a latitude of 24.68 and the north array at 28.75. To obtain conservative constraints, we use the shortest night of the year as the observation period. We assume that in the first night, observations occur from $10^{5}$ s to $8.94\times 10^{4}$ s. In the second night, the observation period extends from $2.93\times 10^{4}$ s to 0 s for the south array and from $2.85\times 10^{4}$ s to 0 s for the north array, corresponding to the entire shortest night of CTA. Based on this, we apply an additional 50% correction to the detected photons. Figure 2 presents the 5-year limit results for CTA before and after this correction.
@@
From Fig. 2, we observe that, without applying the correction, CTA imposes the strongest constraints on PBH burst rate density with a remaining lifetime of $10^{3}$ s. Therefore, to achieve the most stringent limits, CTA should consider setting the observation duration for the same sky region to exceed $10^{3}$ s.
V Conclusions
@@
PBHs, formed in the early universe, are significant candidates for dark matter. Searching for PBHs can enhance our understanding of both the origin and evolution of the universe, as well as the nature of dark matter. This provides the primary motivation for searching local PBH bursts with CTA.
@@
In this paper, we first estimate the number of photons detected by CTA from local PBH bursts, based on detection criteria and background particle rates. We then used the number of detected photons and CTA’s effective area to calculate the maximum detection distance for local PBH bursts. Finally, we estimated the upper limit of the burst rate density, in the situation that no local PBH bursts are detected.
@@
In summary, we showed that CTA can significantly enhance the detection capabilities of PBHs on the parsec scale. We found that the best constraint on the local PBH burst rate density for CTA is $\dot{\rho}<36.26~{}\mathrm{pc}^{-3}\mathrm{yr}^{-1}$. This result is one to two orders of magnitude better than the constraint obtained by LHAASO, as predicted in Ref. Yang et al. (2024), and the current strongest constraints obtained by HAWC. Even when compared to SWGO, another future detector, the best constraint results are several times more stringent. Furthermore, we found that when the observation time for the same sky region is extended to $10^{3}$ s, CTA can achieves the best constraint results. These findings can help optimize detection strategies for future observations to obtain better results. We hope that, in the future, by combining actual observational data from CTA and optimizing observation methods such as considering different search bin sizes, gamma-hadron separation parameters, and selection cuts for different burst durations, as discussed in Ref. Abdo et al. (2015), we can obtain even better constraints.
Acknowledgements.
@@
This work was supported by the National SKA Program of China (grant Nos. 2022SKA0110200, 2022SKA0110203), the National Natural Science Foundation of China (grant Nos. 12473001, 11975072), and the National 111 Project (Grant No. B16009).

### v2

@@ 3 Sensitivities of CTAO in measurements of local burst-rate density
Figure 2: 5-year limit results of CTAO with and without correction.
4 Discussion
@@
Figure 3 presents a comparison between our constraint results and those obtained by other detectors based on observational data [22, 23, 25, 27, 26, 28, 29, 24, 30, 31]. It is evident that CTAO exhibits significantly higher sensitivity in constraining the local PBH burst rate density compared to most previous detectors. The sensitivity of CTAO is only slightly lower than that of LHAASO for bursts with durations shorter than 0.01 s and lower than that of the future SWGO detector for bursts with durations shorter than 1 s. This demonstrates that CTAO holds a distinct advantage in searching for long-duration PBH bursts.
@@
Figure 3: The PBH burst local rate density upper limits of CTAO at 99% confidence level, compared with the results from Whipple [22], CYGNUS [23], H.E.S.S. [24], MAGIC [33], CTAO2015 [32],Tibet Air Shower Array [25], Milagro [26], VERITAS [27], Fermi-LAT [28], HAWC [29], LHAASO [31] and SWGO [30].
@@
To establish a reference for selecting the optimal burst duration when searching for PBH bursts in future observational studies, we analyze the burst durations to which CTAO is most sensitive. This analysis is a key focus of our work. As shown in Table 1, better constraints are likely achieved by targeting PBHs with longer remaining lifetimes. However, since telescopes typically do not observe the same region for extended periods, PBHs with remaining lifetimes significantly longer than the observation time emit more photons than can be detected. Considering this limitation, we apply a rough correction for observations lasting longer than $10^{2}$ s to determine the optimal constraints on PBH bursts for CTAO. Here, the value of $10^{2}$ s refers to the integration window $\Delta t$ used in the PBH burst-search algorithm, rather than the total duration of an observational run. A typical run lasts about $1200$ s for CTA or approximately $28$ minutes for H.E.S.S. [24]. Following previous analyses, such as Ref. [40], we adopt $\Delta t=10^{2}$ s to achieve a balance between signal retention and background accumulation.
We extended the analysis to include remaining lifetimes up to $10^{5}$ s. This extension serves two purposes by identifying the burst duration at which CTAO is most sensitive and evaluating the observational limitations related to PBHs with long remaining lifetimes.
For PBHs with remaining lifetimes of $10^{3}$ s and $10^{4}$ s, we assume that only 50% of the total emitted photons are detected, as the burst duration exceeds the typical integration time. For $\tau=10^{5}$ s, we apply an additional correction due to the limited nighttime visibility. Here, we define time in terms of the PBH remaining lifetime $\tau$, i.e., the time until complete evaporation. Since telescopes cannot observe continuously over such long timescales, we conservatively assume that the observation spans two nights. To maximize realism, we adopt the shortest night of the year as the available window. For the first night, we assume the observation begins when the PBH has $\tau=10^{5}$ s remaining and ends at $\tau=8.94\times 10^{4}$ s, based on the shortest possible dark time. For the second night, the observation window extends from $\tau=2.93\times 10^{4}$ s to $\tau=0$ s for the South array, and from $\tau=2.85\times 10^{4}$ s to $\tau=0$ s for the North array. This corresponds to the maximum usable duration of the shortest night at each site. Combining these two nights covers approximately 11 hours of the 28-hour-long PBH burst. We therefore apply an additional 50% correction to the photon counts for $\tau=10^{5}$ s. The resulting 5-year burst rate density limits for CTAO, before and after applying this correction, are shown in Fig. 2.
@@
From Fig. 2, we observe that, without applying the correction, CTAO imposes the strongest constraints on PBH burst rate density with a remaining lifetime of $10^{3}$ s. Therefore, to achieve the most stringent limits, CTAO should consider setting the observation duration for the same sky region to exceed $10^{3}$ s.
5 Conclusions
@@
PBHs, formed in the early universe, are significant candidates for dark matter. Searching for PBHs can enhance our understanding of both the origin and evolution of the universe, as well as the nature of dark matter. This provides the primary motivation for searching local PBH bursts with CTAO.
@@
In this paper, we first estimate the number of photons detected by CTAO from local PBH bursts, based on detection criteria and background particle rates. We then used the number of detected photons and CTAO’s effective area to calculate the maximum detection distance for local PBH bursts. Finally, we estimated the upper limit of the burst rate density, in the situation that no local PBH bursts are detected.
@@
In summary, we showed that CTAO can significantly enhance the detection capabilities of PBHs on the parsec scale. We found that the best constraint on the local PBH burst rate density for CTAO is $\dot{\rho}<36.26~\mathrm{pc}^{-3}\mathrm{yr}^{-1}$. This result is one to two orders of magnitude better than the constraint obtained by LHAASO, as predicted in Ref. [31], and the current strongest constraints obtained by HAWC. Even when compared to SWGO, another future detector, the best constraint results are several times more stringent. Furthermore, we found that when the observation time for the same sky region is extended to $10^{3}$ s, CTAO can achieves the best constraint results. These findings can help optimize detection strategies for future observations to obtain better results. We hope that, in the future, by combining actual observational data from CTAO and optimizing observation methods such as considering different search bin sizes, gamma-hadron separation parameters, and selection cuts for different burst durations, as discussed in Ref. [26], we can obtain even better constraints.
Acknowledgments
@@
This work was supported by the National SKA Program of China (Grants Nos. 2022SKA0110200 and 2022SKA0110203), the National Natural Science Foundation of China (Grants Nos. 12533001, 12473001, 11975072, 11875102, and 11835009), the China Manned Space Program (Grant No. CMS-CSST-2025-A02), and the 111 Project (Grant No. B16009).